### v1

# Evaluating Multi-Hop Reasoning in Large Language Models: A Chemistry-Centric Case Study

Abstract
In this study, we introduced a new benchmark consisting of a curated dataset and a defined evaluation process to assess the compositional reasoning capabilities of large language models within the chemistry domain. We designed and validated a fully automated pipeline, verified by subject matter experts, to facilitate this task. Our approach integrates OpenAI reasoning models with named entity recognition (NER) systems to extract chemical entities from recent literature, which are then augmented with external knowledge bases to form a comprehensive knowledge graph. By generating multi-hop questions across these graphs, we assess LLM performance in both context-augmented and non-context augmented settings. Our experiments reveal that even state-of-the-art models face significant challenges in multi-hop compositional reasoning. The results reflect the importance of augmenting LLMs with document retrieval, which can have a substantial impact on improving their performance. However, even perfect retrieval accuracy with full context does not eliminate reasoning errors, underscoring the complexity of compositional reasoning. This work not only benchmarks and highlights the limitations of current LLMs but also presents a novel data generation pipeline capable of producing challenging reasoning datasets across various domains. Overall, this research advances our understanding of reasoning in computational linguistics.
1 Introduction
Large Language Models (LLMs) have achieved impressive performance on a wide range of tasks, yet their ability to perform complex, multi-step reasoning remains an ongoing challenge. Techniques such as chain-of-thought (CoT) prompting [1, 2, 3, 4, 5] and structural innovations [6, 7, 8] have enabled notable improvements in reasoning, particularly in mathematics and coding. OpenAI’s o-series [9, 10] was among the first to introduce inference-time scaling of CoT reasoning depth, and subsequent open-source models such as DeepSeek R1 [11, 12] and Qwen QwQ [13] have adopted similar strategies. Notably, these models also leverage reinforcement learning during training to further refine their CoT reasoning, consistently improving performance with increased test-time compute.
To evaluate the reasoning capabilities of LLMs, the community relies on a suite of benchmarks requiring multi-hop reasoning, spanning domains from mathematics [14, 15] and programming [16, 17, 18] to general question answering, including open-book tasks such as HotpotQA [19] and StrategyQA [20]. Recent advances in scaling reasoning have led to improvements in these benchmarks. However, these models remain largely general-purpose. In scientific fields such as chemistry, multi-hop reasoning is essential for integrating interconnected, domain-specific knowledge. Although several multi-hop question answering benchmarks exist, evaluations specific to chemical reasoning are limited [21, 22, 23]. Recent work, including datasets targeting subfields such as reticular chemistry [24], highlights the need for more comprehensive and challenging domain-specific benchmarks.
To address this gap, we propose an automated multi-hop reasoning data generation pipeline that leverages OpenAI’s o3-mini and gpt-4o models. Our pipeline systematically extracts and verifies chemical entities via named entity recognition (NER) and links them to external databases to construct a knowledge graph, from which challenging multi-hop question-answer pairs are generated. Our contributions are as follows:
  1. 1.
    We provide extensive experimental evidence that compositional reasoning in scientific domains remains a significant limitation for current state-of-the-art LLMs.
  2. 2.
    We demonstrate that even retrieval augmentation with perfect context does not guarantee flawless reasoning due to inherent compositional complexities.
  3. 3.
    We introduce an automated knowledge graph and multi-hop reasoning data generation pipeline, leveraging OpenAI models and NER, which can potentially generate unlimited domain-specific datasets.
  4. 4.
    We contribute a challenging benchmark for Multi-hop QA in chemistry
The complete Q&A dataset, together with the evaluation code, is available  [^1] https://github.com/MohammadKhodadad/ChemKGMultiHopQA.
2 Background and Related Works
Multi-hop question answering (QA) has evolved as a key method to evaluate the multi-step reasoning abilities of large language models [25]. Answering multi-hop questions requires integrating multiple pieces of evidence. Traditionally, datasets such as HotpotQA [19], WikiHop, and MedHop [26] were manually or semi-automatically curated through crowd-sourcing and knowledge-base relations. While these approaches yield high-quality, human-validated questions, they are resource-intensive.
More recently, LLMs have been leveraged to generate multi-hop datasets automatically. In many cases, single-hop QA pairs are first generated and later merged using entity linking techniques, a common approach that connects individual entities across questions. For example, the MuSiQue framework [27] fuses two QA pairs by linking a named entity from the first answer to the subsequent question, thereby forming a chain of reasoning. Other methods, such as MultiHop-RAG [28], extend this paradigm by incorporating retrieval-augmented generation (RAG) to paraphrase factual sentences and group them based on shared topics, reflecting the diverse strategies that are emerging in multi-hop QA generation.
Chemistry Domain.
The chemical sciences pose distinct challenges for multi-hop QA due to the need for expert domain knowledge. Only a small fraction of HotpotQA’s questions (roughly 900) are chemistry-focused, limiting both domain relevance and topical currency. Furthermore, they tend to be limited to 2-hop reasoning which limits the difficulty level of the questions. More specialized datasets, such as ChemLitQA [21], provide around 1,000 single-hop and 700 multi-hop question-answer pairs generated using LLMs based on ChemRxiv papers. Although the ChemLitQA-multi dataset includes multi-hop questions, these are typically defined by a single linked entity across all hops. This limitation has led the authors of [21] to emphasize the need for future studies focused on developing more challenging multi-hop QA benchmarks in the field.
Similarly, the chemistry subset of the OlympicArena benchmark [22] provides high-level Olympiad problems that are few in number, and are not explicitly designed for multi-hop reasoning.
Figure 1: An Overview of the knowledge graph generation pipeline.
Knowledge Graph Generation.
Automated knowledge graph construction from unstructured text has seen diverse approaches aimed at effectively capturing entities and their interrelations. Early systems, such as Grapher [29], take an end-to-end approach by first generating nodes with a fine-tuned pretrained language model and then forming edges via sequential generation or classification techniques. Building on these ideas, frameworks like the Extract-Define-Canonicalize (EDC) approach [30] employ a three-phase strategy: they extract relational triplets without a fixed schema, generate natural language definitions for each relation, and then standardize and merge equivalent triplets, sometimes with an additional refinement stage that leverages retrieval techniques.
Complementing these general-purpose methods, dynamic and domain-specific approaches address specialized challenges in KG construction. For instance, KG-MRC [31] models the evolution of entity states in procedural texts by integrating neural reading comprehension with recurrent graph updates to capture temporal changes. In parallel, domain-specific systems such as CEAR [32] incorporate tailored ontologies and specialized knowledge to generate more accurate graphs from the scientific literature. Similarly, Cai et al. [33] demonstrate an iterative, coarse-to-fine refinement process that adapts a broad biomedical knowledge graph to specialized domains like oncology, reducing reliance on manual annotations while preserving essential domain nuances.
3 Methodology
Our methodology consists of three main components: knowledge graph generation, multi-hop question-answer generation, and evaluating state-of-the-art large language models on the question-answering task. The first two components are described in detail in this section, and the last one is explained in the next section.
3.1 Knowledge Graph Generation
We began by constructing a comprehensive knowledge graph from chemical literature. First, using the ChemRxiv API, we collected all ChemRxiv articles with licenses that permitted redistribution. Next, we cleaned the articles using regular expressions to extract their introductions. Focusing on objective and factual information, we extracted each introduction’s first few paragraphs (up to 500 words). Finally, we segmented the extracted text into chunks of up to 128 words, ensuring that no paragraph was split across chunks.
Next, we applied named entity recognition (NER) models to these text chunks to identify chemical entities. In particular, we utilized an NER model [34] that leverages a PubMedBERT architecture [35] fine-tuned on various chemical datasets. To ensure that the extracted entities were specific, verified, and chemically relevant, we utilized OpenAI’s gpt-4o to review and refine the outputs. The same model was also employed to extract relations between these verified entities, forming triplets that capture the interactions and associations present in the text. Additionally, large language models were utilized to extract descriptive features from textual data associated with each entity. To enrich the nodes further during the construction of the knowledge graph, supplementary information from Wikipedia and the PubChem dataset [36] was integrated. Consequently, the finalized knowledge graph comprises nodes representing chemical entities, enhanced by metadata and descriptive annotations from these sources, as well as edges representing the relationships extracted from the textual data. Figure 1 illustrates the procedure followed to generate the knowledge graph.
Figure 2: Overview of the QA generation Pipeline.
3.2 Multi-hop Question-Answer Generation
To generate multi-hop questions, we first sampled paths of varying lengths from the constructed knowledge graph using a randomized breadth-first search (BFS) path sampling algorithm. During path sampling, we ensured that the sources for the edges were distinct, encouraging solutions to integrate information from multiple sources and different parts of the context to answer the questions. Therefore, each path with a length of K involves K+1 entities, coming from K distinct source texts extracted from the original ChemRxiv database.
Adopting a bottom-up approach, we began by sampling paths and generating individual 1-hop questions from each hop. Specifically, For every $(\text{entity}_{1},\text{relation},\text{entity}_{2})$ triplet, we formulated a corresponding question in which $\text{entity}_{1}$ served as the answer, the prompt inquired about the entity that holds the specified relation to $\text{entity}_{2}$. When a question lacked sufficient specificity, we instructed an LLM to enrich it with additional metadata or context from the original text, thereby enhancing clarity and precision.
Model Context Correctness Rate (%) Avg Duration (s) Avg Input Tokens Avg Output Tokens Total Input Tokens (K) Total Output Tokens (K)
Anthropic Claude Sonnet 3.5 V2 40.06 1.54 567 29 550.93 28.69
Anthropic Claude Sonnet 3.5 V2 72.50 1.68 2210 30 2146.11 29.18
Anthropic Claude Sonnet 3.7 44.80 1.61 567 30 550.93 29.35
Anthropic Claude Sonnet 3.7 80.02 1.84 2210 30 2146.11 29.49
Anthropic Claude Sonnet 3.7 (Thinking) 45.73 39.01 583 1777 566.09 1725.79
Anthropic Claude Sonnet 3.7 (Thinking) 84.35 15.35 2228 715 2163.59 694.78
OpenAI GPT-4o-mini 32.34 0.63 204 9 198.60 9.63
OpenAI GPT-4o-mini 62.82 0.71 1628 10 1581.57 10.01
OpenAI GPT-4o 40.27 0.63 204 9 198.60 9.53
OpenAI GPT-4o 68.80 0.71 1628 10 1581.57 9.95
OpenAI o1-mini 41.09 7.78 160 1047 155.88 1017.55
OpenAI o1-mini 71.99 5.68 1609 718 1562.70 697.41
OpenAI o3-mini 47.58 10.84 210 1187 204.43 1153.12
OpenAI o3-mini 80.33 6.12 1634 558 1587.40 542.46
Mistral Large 35.53 0.41 177 13 172.45 13.40
Mistral Large 73.94 0.57 1913 14 1857.70 14.22
Llama 3.3 70B Instruct 32.13 0.33 330 10 320.47 10.56
Llama 3.3 70B Instruct 65.19 0.40 1781 11 1729.91 10.75
Google Gemma 3 27B 32.03 0.89 163 12 158.95 11.94
Google Gemma 3 27B 69.72 1.00 1587 12 1541.55 12.57
DeepSeek R1 44.39 21.06 159 1466 154.40 1423.73
DeepSeek R1 81.98 8.61 1551 573 1506.14 556.55
Qwen QwQ 32B 35.74 68.29 168 2167 163.51 2104.86
Qwen QwQ 32B 79.81 25.18 1665 757 1617.45 735.86
DeepSeek R1 Distill Qwen 32B 34.19 32.04 159 1074 154.70 1043.56
DeepSeek R1 Distill Qwen 32B 79.09 12.11 1633 400 1586.25 389.31
Table 1: Summary of tested models performance in terms of several evaluation metrics for both Contextual and Non-Contextual Setups
These individual questions were then combined into a single multi-hop question using OpenAI’s o3-mini model. Importantly, the final aggregated question was constructed to begin with the last sub-question and chain the entities up to the $\text{entity}_{1}$ of the first relation, ensuring that the final answer corresponds to the answer of the first question. During the verification phase, each one-hop question was first reviewed for clarity, relevance to chemistry, and alignment with the corresponding text that provided the answer. The multi-hop question was then assessed through an additional evaluation step, ensuring that its logical flow effectively led to the final answer. An LLM-based verification process was employed to confirm factual accuracy, answerability based on available context and metadata, and the logical coherence of the sub-questions. Feedback from domain experts was continuously incorporated into the prompts to enhance verification accuracy. To minimize ambiguity, questions that were answered incorrectly by all evaluated models were excluded from the benchmark. Figure 2 illustrates the detailed pipeline of Multi-hop QA generation.
To minimize the impact of writing style and summarization on accuracy evaluation, all questions are designed to have short answers. Answering these questions requires breaking down the main question into smaller sub-questions, finding the answer to each, and combining them to arrive at the final answer. Even with full context available, a correct answer cannot be obtained if the model is not capable of inferring and integrating different pieces of knowledge.
4 Experiments and Results
In our experiments, we evaluated the domain-specific multi-hop question-answering capabilities of a wide variety of state-of-the-art large language models, including both reasoning-focused and general-purpose models. For clarity, throughout this work we refer to models specifically optimized to scale test-time compute as reasoning models. These included open-source and proprietary variants, tested with or without provided context. The summary of tested models and their performance is provided in Table 1.
To access the selected models for evaluation in this experiment, we used different API providers: (i) all tested OpenAI models (gpt-4o, gpt-4o-mini, o1-mini and o3-mini) are accessed via the OpenAI platform; (ii) Amazon Bedrock Platform has been used to access Anthropic Sonnet 3.7 (with and without extended thinking), Anthropic Sonnet 3.5 V2, Mistral Large, DeepSeek R1 and Llama 3.3 70B Instruct; and (iii) Google Gemma 3 27B, Qwen QwQ 32B, and DeepSeek R1 Distill Qwen 32B are accessed via the OpenRouter Platform and operate at bf16 precision. All of these models can perform function-calling tool use, so they were instructed to produce valid JSON outputs to ensure consistency and enable automated validation. All models are evaluated in two settings: with and without provided context. The first scenario reflects performance when the models are paired with an ideal retrieval-augmented generation (RAG) system, while the second scenario relies on the model’s internal memory to answer the questions. After parsing the JSON, we checked whether the output was an exact match to ground truth. If not, OpenAI GPT-4o was instructed to perform a binary assessment; determining whether the answer was correct or not, to calculate the Correctness Rate (%) metric. Our dataset comprises 971 questions spanning 1 to 4 hops (On average 245 questions per hop), generated using the approach described in Section 3.2.
4.1 Models Performance
Figure 3 illustrates the performance of 13 large language models evaluated with respect to correctness rate, cost, and latency in both context-provided and context-not-provided setup. In our performance evaluation, the Llama 3.3 70B Instruct and GPT-4o models achieved the lowest cost and demonstrated notably low latency, but they also registered the lowest correctness rate, making them cost-efficient yet less accurate options. In contrast, Claude Sonnet 3.7 (with extended thinking) achieved the highest correctness rate, albeit at the expense of significantly higher cost and latency. Meanwhile, both Qwen QWEN 32B and Deepseek R1 Distil QWEN 32B maintained a favorable balance between cost and correctness rate when the context is provided – i.e. equipped with a perfect RAG system –, though they incurred above-average latency. Claude Sonnet 3.7 was found to have the highest correctness rate in non-reasoning models.
Figure 3: Performance of selected models based on correctness rate, cost, and latency. The cost axis is log-scaled to emphasize differences. The y-axis represents the percentage of questions answered correctly by each model, while the size of the dots indicates the average latency for each model when answering questions. The top panel displays results for setups where context is provided to the models, while the bottom panel presents results for setups without context. Note that the horizontal axis in both panels has the same range, while the vertical axes ranges differ.
In Context not provided setup, open-source reasoning models (R1 Distill Qwen, QWQ-32B, and R1) tend to have lower performance ranking compared to other models. In contrast, for OpenAI models we observed an opposite trend, which may indicate potentially richer pre-training data. For Claude 3.7, using extended thinking did not lead to better performance compared to the no-thinking setup when the context is not provided to the model; instead, it increased the output tokens and cost. For a comprehensive breakdown of model performance and experimental settings, refer to Table 1, which details metrics such as correctness rate, latency, and token usage. Note that the output details for reasoning models also include the reasoning tokens.
4.2 Comparison with HotpotQA
To demonstrate that large language models, even those designed for reasoning, often struggle with domain-specific multi-hop questions, we created a chemistry-related subset of HotpotQA [19], a well-known general text benchmark primarily sourced from Wikipedia. We sampled chemistry questions by starting from Wikipedia’s Chemistry category, recursively exploring its subcategories (up to three levels), and filtering HotpotQA based on exact title matches. To maintain consistency with our evaluation scheme, we excluded distractors and included only supporting documents as context. Figure 4, illustrates the average performance of all models on each dataset under two conditions: with context provided and without context in the prompt. The results indicate that when context is provided, models achieve similar performance, with our benchmark resulting in marginally lower average performance and reduced variability among the models’ output. In the setup without context, the models found our benchmark more challenging compared to the HotpotQA chemistry subset. This observation may be due to the fact that HotpotQA was exclusively built from Wikipedia, which has been utilized in the pre-training of all evaluated models, while the new benchmark is constructed from more recent ChemRxiv papers enriched with PubChem and Wikipedia.
Figure 4: Comparison of LLMs performance on the chemical subset of HotPotQA with the curated QA dataset from this study. Error bars reflect the standard error of the mean (S.E.M) over evaluated models.
5 Analysis and Ablation
This section presents detailed ablation studies and analyses of the benchmark. First, we examine how context availability and test-time reasoning affect model performance and efficiency. Next, we explore how the number of reasoning hops, an indicator of question difficulty, impacts model correctness and the number of tokens needed to generate an answer.
5.1 Context and Reasoning
In this analysis, we compare the performance of reasoning and non-reasoning models—i.e., models with and without test-time reasoning capabilities, respectively—in two scenarios: with context provided and without context provided. As illustrated in Figure 5-A, providing context significantly improves model performance, nearly doubling the correctness of both reasoning and non-reasoning models. Additionally, reasoning models outperform non-reasoning models in correctly answering questions, benefiting further from their reasoning capabilities in the context-provided setup. As expected and shown in Figure 5B, non-reasoning models are faster than reasoning models. Although these models benefit significantly from context to improve their performance, their latency did not significantly change by providing context to the model. This could be explained by the fact that the number of output tokens in these models is not sensitive to the length of the input prompt and availability of context (see Figure S6). For reasoning models, we observed that the availability of context lowers latency and output token count, likely because available context streamlines the thought process. For further analysis of context and reasoning in these models, refer to Appendix Evaluating Multi-Hop Reasoning in Large Language Models: A Chemistry-Centric Case Study.
Figure 5: Impact of reasoning and context on models’ Correctness Rate (left panel) and latency (right panel). Error bars represent the standard error of the mean for the models in each category.
Figure 6: Analysis of the impact of the number of hops on models performance in Context Provided setup. panel A illustrates the relationship between output token usage (x-axis) and correctness rate (y-axis) for varying numbers of hops. Each point represents a model, while the colored areas indicate distribution estimates for each hop count. Panel B depicts the distribution of correctness rates for non-reasoning models across questions with different numbers of hops, with the dots representing the median of each distribution.
5.2 Impact of the Number of Hops
In this section, we investigate how the number of reasoning hops influences the correctness rate and the output token count. Figure 6 presents the results for the setup with Context Provided. For reasoning models, Figure 6-A illustrates the distribution of answer correctness in relation to the number of generated output tokens. The first observation is that as the number of hops increases, the output token count, which reflects the number of thinking tokens, also increases. However, the answer correctness rate remains relatively constant for multi-hop questions, albeit slightly lower than that observed in single-hop scenarios. For single-hop questions with context provided, we see a decrease in the correctness rate as the output token count increases, indicating a performance trade-off associated with deeper reasoning in simple questions. These trends are not present when context is not provided to the models, as shown in Supplementary Figure S7-A). In non-reasoning models, the output token count is not sensitive to the context or complexity of the question; thus, these models are evaluated solely based on answer correctness. Figure 6-B depicts the distribution of answer correctness rate across the evaluated reasoning models for different numbers of hops. Single-hop questions are answered at a higher correctness rate than multi-hop questions, a trend that is not observed in the setup without context (Supplementary figure S7-B).
6 Conclusion
In this study, we developed a domain-specific multi-hop question-answering (QA) system and evaluated state-of-the-art large language models within the chemistry domain. Our findings reveal that these models struggle with in-domain multi-hop scientific questions, correctly answering fewer than half of the queries when the context is unavailable. Although reasoning fine-tuned models show marginally improved performance, they still face significant challenges. The incorporation of context leads to substantial enhancements, nearly doubling the performance of both reasoning and non-reasoning models. However, even with context, no model, including those fine-tuned for reasoning, achieved a perfect score. Additionally, we contribute to the field by proposing an automated pipeline that integrates advanced named entity recognition with knowledge graph construction to generate intricate multi-hop reasoning tasks, which were utilized for the benchmark. Notably, this potentially domain-agnostic framework can be adapted for various fields by replacing chemistry-specific named entity recognition with suitable alternatives, laying a robust foundation for future research focused on improving reasoning capabilities across diverse specialized domains.
Limitations
Like any research, our study has certain limitations. Specifically, our benchmarking was conducted in two setups: without context and with full context provided. However, in real-world applications, each piece of context is typically collected step by step throughout the reasoning process. This raises the possibility of partial context retrieval, especially for non-reasoning models, which generally rely on a single retrieval round prior to generating answers. Equipping language models with a retrieval-augmented generation pipeline and allowing for multi-step retrieval could introduce a third scenario that better resembles real-world applications. In the past couple of years, advancements in reasoning models have led to various approaches for integrating generative models with RAG systems for multi-step retrieval [28, 37, 38, 39].
Building on these developments, future research could focus on designing and validating an industrial-grade, multi-step retrieval pipeline tailored to chemical text. This enhanced system will support incremental context acquisition and structured reasoning over lengthy or loosely organized documents, thereby enabling accurate answers to more challenging chemical queries.
Acknowledgements
The author(s) gratefully acknowledge the financial support provided by MITACS under funding number IT32409 for the research leading to the publication of this article. We also thank Adam Wojciech Bartwiki for project management support, Tobias Roth for his essential contributions to the infrastructure, and Stephen Dokas for his invaluable recommendations in the chemistry domain.
Appendix
Appendix S1 Detailed Performance Based on Context Availability
Figure S1 shows that model performance is strongly influenced by whether context is provided in the input. In particular, Claude 3.7 with extended thinking achieved a correctness rate of 84% with context, whereas o3-mini recorded the highest correctness rate (48%) when the context was absent. Note that o3-mini was primarily used to generate the questions, which may have introduced a slight bias resulting in its minor improvement in correctness. Figures S2 and S3 illustrate the token usage and latency of the models, respectively.
Figure S1: Correctness rate of models with respect to context.
Figure S2: Token usage of models with respect to context
Figure S3: Latency with respect to context
S1.1 Reasoning without Provided Context
Figure S4 illustrates the performance of various models when no context is provided. As shown, o3-mini achieves the highest correctness rate, although this comes at a higher cost. It is important to note that o3-mini was primarily used for question generation, which may provide it with a slight advantage. In contrast, Llama3-3 70B Instruct records the lowest cost and lowest correctness rate while also exhibiting low latency. Meanwhile, Claude Sonnet 3.7 (without thinking tokens) balances correctness and cost, with acceptably low latency. This model also exhibited the highest correctness rate among non-reasoning models.
Figure S4: Performance of selected models across correctness rate, cost, and latency. The cost axis is log-scaled to highlight differences. The y-axis represents the average performance of models in providing the correct answer, and the dot size indicates model latency.
Appendix S2 Performance of models on Chemistry Subset of HotpotQA
Table S1 shows the details of each model performance, latency, and tokens used in both setups of context provided and not provided for the chemistry subset of HotpotQA [19] questions.
Model Context Correctness Rate (%) Avg Duration (s) Avg Input Tokens Avg Output Tokens Total Input Tokens (K) Total Output Tokens (K)
Anthropic Claude Sonnet 3.5 V2 53.27 1.26 517 29 507.47 29.20
Anthropic Claude Sonnet 3.5 V2 84.80 1.32 618 29 605.73 28.85
Anthropic Claude Sonnet 3.7 58.27 1.87 517 29 507.47 28.49
Anthropic Claude Sonnet 3.7 86.22 1.94 618 29 605.73 28.71
Anthropic Claude Sonnet 3.7 (Thinking) 65.31 17.54 539 726 528.48 711.49
Anthropic Claude Sonnet 3.7 (Thinking) 87.14 10.10 640 390 627.29 382.77
OpenAI GPT-4o-mini 45.61 0.43 170 7 166.96 7.72
OpenAI GPT-4o-mini 80.31 0.50 257 8 252.39 8.00
OpenAI GPT-4o 55.10 0.62 170 8 166.96 8.76
OpenAI GPT-4o 81.33 0.63 257 8 252.39 8.72
OpenAI o1-mini 50.82 4.82 127 719 124.82 705.00
OpenAI o1-mini 85.51 3.41 217 426 213.30 418.09
OpenAI o3-mini 59.69 9.65 166 977 163.04 957.92
OpenAI o3-mini 87.65 3.74 253 247 248.47 242.51
Mistral Large 4.59 0.49 198 18 194.44 17.86
Mistral Large 0.92 0.36 303 12 297.72 12.32
Llama 3.3 70B Instruct 44.49 0.32 284 9 278.92 9.04
Llama 3.3 70B Instruct 79.29 0.29 373 9 365.78 9.03
Google Gemma 3 27B 40.51 0.77 127 10 124.58 9.86
Google Gemma 3 27B 79.80 0.82 218 11 214.04 11.26
DeepSeek R1 59.08 8.93 125 612 122.76 600.15
DeepSeek R1 85.10 5.12 212 358 208.25 351.01
Qwen QwQ 32B 51.94 27.91 126 865 124.45 848.04
Qwen QwQ 32B 88.37 11.01 219 412 215.34 403.94
DeepSeek R1 Distill Qwen 32B 46.63 16.20 119 565 117.10 553.77
DeepSeek R1 Distill Qwen 32B 86.84 7.98 208 287 204.53 282.17
Table S1: Summary of tested models performance on HotpotQA chemistry subset in terms of several evaluation metrics for both Contextual and Non-Contextual Setups
Appendix S3 A Multi-Hop QA Generation Example
Figure S5 illustrates a typical multi-hop QA example derived from our knowledge-graph-based methodology. The context is drawn from chemical literature discussing the use of carbon dioxide as a renewable feedstock for formic acid, which then serves as a non-gaseous CO surrogate in carbonylation reactions. By chaining these facts together, our approach constructs a question that requires integrating multiple pieces of information to arrive at the correct answer. This demonstrates how multi-hop reasoning, guided by entity relations and supplemented with descriptive metadata, enables more complex question generation and evaluation of large language models. Additionally, Figure 2 shows the step-by-step process of deriving multi-hop questions from a knowledge graph, illustrating how entities, relations, and descriptive metadata are combined to construct more complex queries.
Context:
Carbonylation reactions constitute a potent tool to manufacture carboxylic acids and their derivatives both in industry and academic organic synthesis. In general, carbonylation requires the use of toxic carbon monoxide, which thus usually demands certified high-pressure reaction vessels. Therefore, developing non-gaseous CO surrogate for conducting safe and facile-operation carbonylation is an important and ongoing research topic. Among these established CO surrogates, formic acid is one kind of versatile atom.
The utilization of carbon dioxide as a C1 feedstock for the generation of industrially relevant chemicals is also an interesting approach. CO2 is an attractive renewable C1 source, which can lead to formic acid. Those approaches would not only reduce carbon dioxide emissions through carbon capture but also compensate sequestration costs by producing chemicals in global demand. Question:
What is the process that uses a substance, produced from carbon dioxide and known as the simplest carboxylic acid with antibacterial and preservative properties, as a non-gaseous surrogate to safely form carboxylic acids and their derivatives under mild conditions?
Answer: carbonylation reactions Sentence-level supporting facts:
1) formic acid is the simplest carboxylic acid with antibacterial and preservative properties.
2) formic acid can be produced from carbon dioxide. 3) formic acid can act as a non-gaseous CO surrogate. 4) carbonylation reactions safely produce carboxylic acids under mild conditions using formic acid as a CO surrogate. Path (multi-hop chain of reasoning):
carbon dioxide
$\rightarrow$ formic acid $\rightarrow$ carbonylation reactions
Figure S5: An example of a multi-hop question-answer.
Appendix S4 Impact of Context and Reasoning on output tokens count
Figure S6 illustrates the impact of context availability on the average number of output tokens generated by reasoning and non-reasoning models when answering questions. Non-reasoning models produce a similar number of tokens, as they do not engage in test-time reasoning. In contrast, for reasoning models, the number of tokens generated to answer questions decreases with the availability of context, indicating a potential requirement of less thinking when the context is available.
Figure S6: Visualization of the token usage distribution based on input context availability and model reasoning capability. The y-axis is log-scaled.
S4.1 Performance Analysis Based on Number of Hops
Figure S7 illustrates how the number of hops affects performance in the absence of context. The data reveals that as the number of hops increases, the correctness rate declines while the number of output tokens rises. Additionally, both Figure LABEL:fig:hops_context and Figure S7 show a negative correlation between token count and correctness when only one hop is used. This may suggest that overanalyzing simpler questions could lead to errors in the final answer. We visualized the overall performance of all models in Figure S8 and  S9 to analyze how the correctness rate varies with the number of reasoning hops. In addition, token usage and latency metrics were separately depicted in Figures S10,  S11,  S12 and S13, respectively, to provide a more detailed view of the efficiency and resource consumption as the reasoning depth increases.
Figure S7: Analysis of the impact of the number of hops on model performance in the absence of context. A) Scatter and KDE plots showing the relationship between the number of hops, correctness rate, and output tokens. B) Distribution of non-reasoning models’ performance across different numbers of hops.
The following figures illustrate each model’s performance, latency, and token usage for question clusters that require different numbers of hops to answer.
Figure S8: Overall performance of models as a function of the number of hops When context is provided.
Figure S9: Overall performance of models as a function of the number of hops when context is not provided.
Figure S10: Token usage across different numbers of hops when context is provided.
Figure S11: Token usage across different numbers of hops when context is not provided.
Figure S12: Latency measurements across different numbers of hops when context is provided.
Figure S13: Latency measurements across different numbers of hops when context is not provided.